# TabMGP: Martingale Posterior with TabPFN

### v1

@@ Abstract
Bayesian inference provides principled uncertainty quantification but is often limited by challenges of prior elicitation, likelihood misspecification, and computational burden. The martingale posterior (MGP, Fong et al., 2023) offers an alternative, replacing prior–likelihood elicitation with a predictive rule — namely, a sequence of one-step-ahead predictive distributions — for forward data generation. The utility of MGPs depends on the choice of predictive rule, yet the literature has offered few compelling examples. Foundation transformers are well-suited here, as their autoregressive generation mirrors this forward simulation and their general-purpose design enables rich predictive modeling. We introduce TabMGP, an MGP built on TabPFN, a transformer foundation model that is currently state-of-the-art for tabular data. TabMGP produces credible sets with near-nominal coverage and often outperforms both existing MGP constructions and standard Bayes.
@@ 1 Introduction
Bayesian inference provides a principled framework for uncertainty quantification, but its classical formulation hinges on the specification of a prior–likelihood pair. This requirement can be burdensome in practice. Priors are often placed on uninterpretable parameters, and “uninformative” defaults can dominate inference in data-scarce or high-dimensional regimes (Zhang et al., 2022). The problem is further compounded by the fact that the likelihood itself is often misspecified in modern applications. These limitations have motivated a wave of generalisations – ranging from generalised Bayes (Bissiri et al., 2016) to optimisation-centric Bayes (see e.g. Knoblauch et al., 2022; Wild et al., 2023; Shen et al., 2025) that modernise parts of the Bayesian recipe. Collectively, they mark the emergence of a post-Bayesian paradigm in statistical inference.
The martingale posterior represents one direction of this expansion. Originating in the Bayesian predictive literature (Fortini and Petrone, 2020; Berti et al., 2021; Fong et al., 2023), the martingale posterior shifts the focus from prior specification to prediction. Instead of eliciting a prior over parameters, the analyst provides a predictive rule, defined as a sequence of one-step-ahead predictive distributions. When such predictive rules satisfy certain technical conditions — namely, the martingale property — the resulting posterior is guaranteed to exist and be well-behaved (formal definition in Section 2). Posterior inference is then obtained by forward sampling: starting from the observed dataset, one repeatedly draws the next observation from the predictive rule conditional on the data so far, append it to the dataset, and continue.
The Bayesian predictive approach is particularly attractive in the modern era of powerful black-box predictive models, especially foundation transformers (Bommasani et al., 2022). In fact, the forward sampling mechanism described above will be familiar from large language models (LLMs): the predictive rule plays the role of next-token prediction, and forward sampling mirrors autoregressive generation, where at each step a token is sampled from the conditional distribution given the context, appended to the context, and the process is repeated. In this sense, transformers are naturally aligned with the forward-sampling procedure that underpins the martingale posterior. Furthermore, pretrained foundation models are explicitly designed to deliver accurate predictions across broad domains without bespoke modelling, making them natural candidates for predictive rules.
@@
Yet, despite this alignment, the martingale posterior literature has focused almost exclusively on handcrafted predictive rules that satisfy exact martingale properties (e.g., Fong et al., 2023; Ghalebikesabi et al., 2023; Huk et al., 2024); see Section 4 for a review. These approaches are elegant, but they require problem-specific design choices and often impose strong conditions. In addition, the emphasis on exact martingale structure risks overstating its importance: the martingale property is sufficient to guarantee the existence of a limiting law and a well-defined martingale posterior, but it is not clear that it is necessary for delivering meaningful expressions of uncertainty about unknown parameters.
@@
This motivates our central question: can modern predictive rules be used as the predictive rule in martingale posteriors? We answer in the affirmative by introducing TabMGP, the first martingale posterior powered by a large pretrained foundation transformer. Specifically, we instantiate the predictive rule with TabPFN (Hollmann et al., 2022; 2025), a transformer trained on a broad range of synthetic tabular data that now achieves state-of-the-art results for tabular classification and regression tasks. This integration yields several benefits:
  • Off-the-shelf adaptivity: TabMGP requires no bespoke design of predictive rules or task-specific fitting.
  • Scalability and generality: TabMGP applies broadly across regression and classification tasks, requiring only the specification of a loss functional.
Our results show that TabMGP consistently provides reliable uncertainty quantification, often matching or surpassing both handcrafted martingale posteriors and classical Bayesian inference. This establishes a practical new method for the modern Bayesian’s toolkit and open new avenues and questions for broadening the martingale posterior framework.
@@ 2 Martingale posterior
The martingale posterior (Fong et al., 2023) is a recently proposed statistical inference method within the predictive Bayesian inference framework (see Fortini and Petrone, 2025, for a review). We largely follow the notation of Fortini and Petrone (2025) in our review of the martingale posterior framework below.
Notation.
Let $(Z_{i})_{i\geq 1}$ be random variables taking values in a space $\mathcal{Z}$}. For any probability law $\mathbb{P}$} for the sequence $(Z_{i})_{i\geq 1}\in{\mathcal{Z}}^{\infty}$, define the predictive rule as the sequence of predictive distribution $P_{0}(\cdot)\coloneq{\mathbb{P}}(Z_{1}\in\cdot)$ and, for $i\geq 1$,
@@
We write $P_{i}(\cdot\mid z_{1:i})$ for the realisation given $z_{1:i}$. We also write $\mathbb{P}$}-a.s. for “with $\mathbb{P}$}-probability one”. We use $p$ to denote the probability density or mass function of $P$; for example, $p_{i}$ is the density or mass function of $P_{i}$.
@@
    A functional of interest $\theta:{\mathcal{P}}({\mathcal{Z}})\to\Theta\subseteq\mathbb{R}^{p}$, where ${\mathcal{P}}({\mathcal{Z}})$ denotes the set of all probability distributions $F$ on $\mathcal{Z}$}.
@@
For the martingale posterior distribution to be well-defined, the sequence of predictive distributions $(P_{i})_{i\geq 0}$ must converge $\mathbb{P}$}-a.s. to a random probability measure $F_{\infty}$, which represents the limiting empirical distribution of $(Z_{i})_{i\geq 1}$. The probability law of $F_{\infty}$ can be regarded as a prior distribution and the probability law of $F_{\infty}\mid z_{1:n}$ a posterior distribution. This posterior law of $F_{\infty}\mid z_{1:n}$ then induces a posterior distribution on functionals of $F_{\infty}$ which is termed the martingale posterior.
Definition.
Let $z_{1:n}\in{\mathcal{Z}}^{n}$ denote the observed data and $\Pi(F_{\infty}\mid z_{1:n})$ the posterior law of $F_{\infty}$ given $z_{1:n}$. The martingale posterior distribution is defined as the posterior law of $\theta(F_{\infty})$:
@@
for all measurable sets $A\subseteq\Theta$. The classical Bayesian posterior can be recovered from (2) as a special case by using a Bayesian posterior predictive distribution (PPD) as the predictive rule and the corresponding negative log-likelihood as the loss function; see Appendix B for details.
Sufficient conditions
A sufficient, but not necessary, condition for the existence of $F_{\infty}$ and hence for the martingale posterior distribution to be well-defined, is that the sequence $(P_{i})_{i\geq 0}$ is a martingale, or equivalently, that the sequence $(Z_{i})_{i\geq 1}$ is conditionally identically distributed (c.i.d, Berti et al., 2004). The martingale property stipulates that $(P_{i})_{i\geq 0}$ satisfies, for every $i\geq 0$ and every measurable set $A$:
$\mathbb{E}(P_{i+1}(A)\mid Z_{1:i})=P_{i}(A).$ (3)
Relaxations of this sufficient condition remain an active area of research (see, e.g., Battiston and Cappello, 2025). As noted in the introduction, the term martingale posterior is somewhat of a misnomer. Any predictive rule with a well-defined $F_{\infty}$ limit induces a posterior distribution as in (2). In this paper, we use the term martingale posterior in a broad sense: even predictive rules that violate the martingale property may still induce valid “martingale” posteriors.
Computation.
The martingale posterior (2) generally does not admit a closed-form expression and must be approximated by a finite martingale posterior, which is the posterior law of $\theta(F_{N})$ where $F_{N}$ is the empirical distribution of $Z_{1:N}$. A sample from the finite martingale posterior can then be obtained as follows:
  1. 1.
    Starting from $z_{1:n}$, sequentially forward sample future observations $Z_{n+1:N}$ for some large $N$:
    $Z_{i+1}\sim P_{i},\quad i=n,\dots,N-1;$
  2. 2.
    Form $F_{N}$ and compute $\theta(F_{N})$.
This procedure for generating a sample from the martingale posterior is known as predictive resampling (Fong et al., 2023). To approximate the (finite) martingale posterior, we repeat predictive resampling to obtain $L$ samples $\{\theta^{(l)}\}_{l=1}^{L}$. The computation of these samples is embarrassingly parallelisable, as predictive resampling can be carried out independently for each sample.
@@ 3.1 Proposed methodology
We propose instantiating the martingale posterior with the predictive rule induced by TabPFN through in-context learning. TabPFN is a prior-data fitted network (Müller et al., 2022), explicitly trained to approximate a Bayesian posterior predictive distribution (PPD) by sampling training data from exchangeable sequences defined by prior–likelihood pairs; see Appendix A for a review. We term the resulting martingale posterior TabMGP.
Several aspects of TabPFN directly support its role as a predictive rule in martingale posteriors:
  1. 1.
    In-context learning. At inference, TabPFN consumes both the observed context $z_{1:i}$ and a new query $x_{i+1}$, and returns a one-step-ahead predictive distribution
    $P_{i}(\cdot\mid x_{i+1})\coloneq{\mathbb{P}}(Y_{i+1}\in\cdot\mid x_{i+1},z_{1:i}),$
    without retraining or fine-tuning. This property allows TabPFN to adapt to new datasets on the fly. We note that TabPFN can handles both classification and regression tasks: in classification it outputs a deterministic probability vector for $y_{i+1}$, while in regression it effectively performs multi-class classification by binning the response variable $y$.
  2. 2.
    Row-permutation invariance. Predictions in TabPFN depend only on the multiset $\{z_{1},\dots,z_{i}\}$, not on the order of its elements. This invariance is built into the TabPFN architecture, whereas many existing martingale posteriors (notably copula-based constructions, see Section 4) must enforce it by averaging predictive distributions over permutations of the feature order — a process that is computationally expensive.
  3. 3.
    Bayesian alignment. As a prior-data fitted network, TabPFN is explicitly trained to approximate Bayesian posterior predictive distributions (PPDs). In the martingale posterior framework, the Bayesian PPD is in some sense the ideal predictive rule: if the predictive rule coincides with the true Bayesian PPD, then the resulting martingale posterior recovers the classical Bayesian posterior (see Appendix B). Training TabPFN on synthetic data generated from a wide range of prior–likelihood pairs encourages it to approximate this ideal behaviour across diverse tasks. This makes TabPFN’s predictive distributions a more principled choice for constructing martingale posteriors.
Why not other transformers?
@@
Large language models and other pretrained transformers often exhibit in-context learning: given a context $z_{1:i}$ and query $x_{i+1}$, they can produce a prediction for $y_{i+1}$. In this sense, they share the same basic mechanism as TabPFN. However, they lack the two further properties that make TabPFN especially well-suited as a predictive rule for martingale posteriors.
@@
First, they are not row-permutation invariant. Because such models rely on positional embeddings, their predictions change with the ordering of training examples, unless explicit averaging is performed across permutations. Second, they are not trained to approximate Bayesian posterior predictive distributions: their training objectives are tailored to next-token prediction in text rather than calibrated probabilistic prediction for supervised tabular tasks. These differences mean that, although generic transformers can perform in-context learning, their predictive distributions are not naturally aligned with the martingale posterior framework. By contrast, TabPFN combines in-context learning with built-in permutation invariance and Bayesian alignment, making it a more principled choice.
@@
Require : Observed data $z_{1:n}$, loss function $\ell(z,\theta)$, forward sampling depth NN, TabPFN $P_{i}(\cdot\mid x_{i+1})$, number of posterior samples $L$
@@
      Sample $y_{i+1}^{(l)}\leftarrow{P}_{i}(\cdot\mid x_{i+1}^{(l)})$;
@@
Validity of TabMGP as a martingale posterior.
As noted in Section 2, we use the term “martingale posterior” in a broad sense: any predictive rule that induces a well-defined limiting distribution $F_{\infty}$ gives rise to a posterior of the form (2), even if the rule does not satisfy the martingale or the almost conditionally identically distributed (a.c.i.d.) conditions (Battiston and Cappello, 2025), which are commonly used as sufficient guarantees. TabPFN does not meet these conditions (see Appendix C), but this does not preclude TabMGP from being a valid instance of the martingale posterior. Establishing more general criteria for the existence of $F_{\infty}$ remains an open theoretical question. In the meantime, we assess validity empirically through diagnostics such as trace plots for the convergence of $\theta(F_{N})$ (Section 5) and concentration properties (Appendix C). Our empirical results indicate that TabMGP does converge and concentrate appropriately in practice.
Areas for improvement.
TabMGP inherits several architectural limitations of TabPFN. One issue is that continuous responses are discretised through binning, effectively reducing regression to classification. While this enables TabPFN to reuse its classification machinery, it introduces approximation error that can distort uncertainty quantification. Another limitation is the lack of invariance to column permutations: predictions may depend on the arbitrary ordering of features, and TabPFN mitigates this by averaging over multiple random permutations at inference time. Although effective, a truly column-permutation-invariant architecture would be more principled and efficient. A further limitation concerns TabPFN’s training objective: it optimises a final prediction loss rather than the autoregressive loss standard in transformer architectures. As a result, forward sampling requires repeated calls to the model, making TabMGP computationally heavy. If TabPFN were trained autoregressively, forward sampling could be implemented in a single network pass, greatly reducing cost.

### v2

@@ Abstract
Bayesian inference provides principled uncertainty quantification but is often limited by challenges of prior and likelihood elicitation. The martingale posterior (MGP) (Fong et al., 2023) offers an alternative by replacing these requirements with a predictive rule. Additionally MGP focuses inference on parameters defined through a loss function. This framework is especially resonant in the era of foundation transformers; practitioners increasingly leverage models like TabPFN for their state-of-the-art capabilities, yet often require epistemic uncertainty for a scientific estimand $\theta$ that need not parameterise the model’s implicit latent model. The MGP provides the mechanism to recover these posterior distributions. We introduce TabMGP, an MGP built on TabPFN for tabular data. TabMGP produces credible sets with near-nominal coverage and often outperforms both handcrafted MGP constructions and standard Bayesian baselines.
@@ 1 Introduction
[…]
Figure 1: TabMGP for obtaining posterior samples of $\theta(F_{\infty})\mid z_{1:n}$, where each $z=(x,y)$ represents a covariate-response pair. We perform forward sampling with TabPFN to generate $L$ independent continuations of the observed dataset $z_{1:n}$. Since TabPFN does not model covariates, the forward sampling of $x_{i}$ is performed via a separate process, while TabPFN provides the conditional response $y_{i}\sim\mathrm{TabPFN}(\cdot\mid x_{i},z_{1:i-1})$. At the termination of each rollout $l$, we form the empirical measure $F_{N}^{(l)}=\frac{1}{N}\left(\sum_{i=1}^{n}\delta_{z_{i}}+\sum_{i=n+1}^{N}\delta_{z_{i}^{(l)}}\right)$ and collect $\theta(F_{N}^{(l)})$ as one approximate posterior sample.
@@
Classical Bayesian inference provides a principled framework for uncertainty quantification, but its formulation hinges on the specification of a prior–likelihood pair. This requirement can be burdensome in practice. Priors are often placed on uninterpretable parameters, and “uninformative” defaults can dominate inference in data-scarce or high-dimensional regimes (Zhang et al., 2022). The problem is further compounded by the frequent misspecification of the likelihood in modern applications; even when a model is settled upon, approximating the resulting posterior—particularly in high dimensions—remains a formidable computational challenge.
These limitations have catalysed a wave of generalisations—ranging from Bayesian predictive inference (BPI) (Fortini and Petrone, 2025) to generalised Bayes (GB) (Bissiri et al., 2016) and optimisation-centric Bayes (see e.g. Knoblauch et al., 2022; Wild et al., 2023; Shen et al., 2025)—that collectively mark the emergence of a post-Bayesian paradigm. Within this paradigm, BPI shifts the focus from prior and likelihood elicitation to the sequence of predictions themselves. Specifically, the analyst provides a predictive rule, defined as a sequence of one-step-ahead predictive distributions $(P_{i})_{i\geq 0}$. BPI is post-Bayesian in the sense that the predictive rule need not correspond to a standard Bayesian posterior prediction distribution (PPD). Instead, the framework requires only that the sequence $(P_{i})_{i\geq 0}$ converges weakly to random limiting measure $F_{\infty}$ almost surely.
Then, we may ask for the posterior distribution of $F_{\infty}$ given data $z_{1:n}$ for some measurable set $A$. Here, $F_{\infty}$ represents the latent Bayesian model internalised by the predictive rule—a state of knowledge that remains implicit but can be queried through simulation. Specifically, we may obtain posterior samples of $F_{\infty}\mid z_{1:n}$ through a process of forward sampling: starting from the observed dataset $z_{1:n}$, one repeatedly draws the next observation from the predictive rule conditional on the data so far, append it to the dataset, and continue. In practice, the forward sampling process is terminated at some large $N$ such that $F_{N}$ serves as an approximate posterior sample of $F_{\infty}\mid z_{1:n}$.
Complementary to BPI is GB, a framework that argues we can update a prior to a posterior distribution over parameters that are connected to observations through a loss function rather than a likelihood. This allows for significant decision-theoretic flexibility: the object of interest $\theta$ is generally defined as the minimiser of an expected loss and need not index the underlying Bayesian model. While GB allows for decision-theoretic flexibility, it still requires the specification of an explicit prior.
The martingale posterior (MGP) emerges as a powerful synthesis of these directions, one that is especially resonant in the era of foundation transformers (Bommasani et al., 2022), which have recently begun to dominate tabular data tasks—a domain long considered the holdout of traditional machine learning, i.e., non-deep learning, methods. Practitioners confronted with a new tabular dataset will increasingly turn to pretrained models like TabPFN (Hollmann et al., 2022) rather than design a bespoke likelihood–prior pair and navigate the complexities of Markov chain Monte Carlo or variational inference. For these users, the parameters of scientific interest are effectively decoupled from the internals of the transformer, necessitating a framework that can recover the posterior distribution of a functional $\theta(F_{\infty})$ directly from the transformer’s predictive outputs. The MGP provides this mechanism: it treats the transformer as a predictive engine for the latent model $F_{\infty}$, while the GB loss function specifies the functional $\theta$ to be extracted.
The forward sampling mechanism we described above for $F_{\infty}\mid z_{1:n}$ provides the computational means for this recovery. The procedure remains the same except, at termination, we report $\theta(F_{N})$ as an approximate posterior sample from $\theta(F_{\infty})\mid z_{1:n}$. Note the sampling is well suited to transformer architectures: next-token prediction is the predictive rule, and forward sampling mirrors autoregressive generation, where each token is sampled from the conditional distribution given the context, appended to the context, and the process is repeated. This procedure is visualised in Figure 1.
@@
Yet, despite this alignment, the martingale posterior literature has focused almost exclusively on handcrafted predictive rules that satisfy exact martingale properties (e.g., Fong et al., 2023; Ghalebikesabi et al., 2023; Huk et al., 2024); see Section 4 for a review. These approaches are theoretically rigorous, but they require problem-specific design choices to tune the predictive rule for the dataset at hand. Moreover, the emphasis on exact martingale structure overstates the importance of this condition: while the martingale property is a sufficient condition to guarantee the existence of a limiting law $F_{\infty}$, it is not necessary for delivering epistemic uncertainty for $\theta(F_{\infty})$. We argue that the reliance on this restrictive sufficient condition has acted as impediment to the growing field of MGP, preventing the application of MGPs to high-capacity models.
@@
This motivates our central question: can modern foundation transformers be used as the predictive rule in martingale posteriors? We answer in the affirmative by introducing TabMGP, the first martingale posterior powered by a large pretrained foundation transformer. Specifically, we instantiate the predictive rule with TabPFN (Hollmann et al., 2022; 2025), a transformer trained on a broad range of synthetic tabular data that now achieves state-of-the-art results for tabular classification and regression tasks.
Our results show that TabMGP consistently provides reliable uncertainty quantification, often matching or surpassing both handcrafted martingale posteriors and classical Bayesian inference. These findings demonstrate that the MGP framework is considerably more robust than initially theorised: even when high-capacity predictive rules like TabPFN do not strictly satisfy formal sufficient conditions such as the martingale property, the resulting posteriors remain empirically stable, providing superior frequentist coverage and tighter credible sets than standard baselines. By establishing a new method for the modern Bayesian toolkit, this work opens new avenues for broadening the MGP framework to the era of foundation models.
@@ 2 Martingale posterior
The martingale posterior (Fong et al., 2023) is a recently proposed statistical inference method within the BPI framework (see Fortini and Petrone, 2025, for a review). We follow the notation of Fortini and Petrone (2025) in our discussion.
Notation. Let $(Z_{i})_{i\geq 1}$ be random variables taking values in a space $\mathcal{Z}$}. For any probability law $\mathbb{P}$} for the sequence $(Z_{i})_{i\geq 1}\in{\mathcal{Z}}^{\infty}$, define the predictive rule as the sequence of predictive distribution $P_{0}(\cdot)\coloneq{\mathbb{P}}(Z_{1}\in\cdot)$ and, for $i\geq 1$,
@@
We write $P_{i}(\cdot\mid z_{1:i})$ for the realisation given $z_{1:i}$. We also write $\mathbb{P}$}-a.s. for “with $\mathbb{P}$}-probability one”. We use $p_{i}$ to denote the density or mass function of $P_{i}$.
@@
    A functional of interest $\theta:{\mathcal{P}}({\mathcal{Z}})\to\Theta\subseteq\mathbb{R}^{p}$, where ${\mathcal{P}}({\mathcal{Z}})$ denotes the set of all distributions $F$ on $\mathcal{Z}$}.
@@
For the martingale posterior distribution to be well-defined, the sequence of predictive distributions $(P_{i})_{i\geq 0}$ must converge $\mathbb{P}$}-a.s. to a random probability measure $F_{\infty}$, which represents the limiting empirical distribution of $(Z_{i})_{i\geq 1}$. Here, $F_{\infty}$ is analogous to the statistical model in classical Bayes, and the probability law of $F_{\infty}$ and $F_{\infty}\mid z_{1:n}$ are the prior and posterior respectively. This posterior law of $F_{\infty}\mid z_{1:n}$ then induces a posterior distribution on functionals of $F_{\infty}$ which is termed the martingale posterior.
Definition. Let $z_{1:n}\in{\mathcal{Z}}^{n}$ denote the observed data and $\Pi(F_{\infty}\mid z_{1:n})$ the posterior law of $F_{\infty}$ given $z_{1:n}$. The martingale posterior distribution is defined as the posterior law of $\theta(F_{\infty})$:
@@
for all measurable sets $A\subseteq\Theta$. The classical Bayesian posterior can be recovered from (2) as a special case by using a Bayesian PPD as the predictive rule and the corresponding negative log-likelihood as the loss function (Appendix B).
Sufficient conditions A sufficient, but not necessary, condition for the existence of $F_{\infty}$ and hence for the martingale posterior distribution to be well-defined, is that the sequence $(P_{i})_{i\geq 0}$ is a martingale, or equivalently, that the sequence $(Z_{i})_{i\geq 1}$ is conditionally identically distributed (c.i.d.) (Berti et al., 2004). The martingale property stipulates that $(P_{i})_{i\geq 0}$ satisfies, for every $i\geq 0$ and every measurable set $A$:
$\mathbb{E}[P_{i+1}(A)\mid Z_{1:i}]=P_{i}(A).$ (3)
Relaxations of this sufficient condition remain an active area of research (see, e.g., Battiston and Cappello, 2025). As noted in the introduction, the term “martingale posterior” is somewhat of a misnomer. Any predictive rule with a well-defined $F_{\infty}$ limit induces a posterior distribution as in (2). In this paper, we use the term in a broad sense: even predictive rules that violate the martingale property may still induce valid “martingale” posteriors.
Computation. The martingale posterior (2) generally does not admit a closed-form expression and is approximated by the finite martingale posterior. This is the posterior law of $\theta(F_{N})$, where $F_{N}$ is the empirical distribution of $Z_{1:N}$. A draw from $\theta(F_{N})$ is obtained by forward sampling a simulation rollout $Z_{n+1:N}$ from the predictive rule:
$Z_{i+1}\sim P_{i},\quad i=n,\dots,N-1,$
then forming $F_{N}$ and computing $\theta(F_{N})$ (Fong et al., 2023). Repeating this produces $L$ samples $\{\theta^{(l)}\}_{l=1}^{L}$. These rollouts are independent and thus embarrassingly parallel.
@@ 3.1 Proposed methodology
We propose instantiating the martingale posterior with the predictive rule induced by TabPFN through in-context learning. TabPFN is a prior-data fitted network (Müller et al., 2022), explicitly trained to approximate a Bayesian PPD by sampling training data from exchangeable sequences defined by likelihood–prior pairs; see Appendix A for a review. We term the resulting martingale posterior TabMGP.
Several distinctive features of TabPFN facilitate its integration as a predictive rule within the martingale posterior framework. Through in-context learning, TabPFN consumes the context alongside a new query to return a one-step-ahead predictive distribution without the need for retraining or fine-tuning. This allows the model to adapt to novel datasets instantaneously, handling classification via deterministic probability vectors and regression through a multi-class classification approach that bins the response variable.
The architecture further benefits from native row-permutation invariance, ensuring that predictions depend solely on the multiset rather than the sequence order. This built-in property provides a significant computational advantage over many existing martingale posteriors, such as copula-based constructions (Fong et al., 2023), which must manually enforce invariance by averaging over permutations. Finally, as a prior-data fitted network, TabPFN is designed to be approximately Bayesian, having been trained to approximate Bayesian PPD. In the martingale posterior framework, the Bayesian PPD serves as the ideal predictive rule; when they coincide, the martingale posterior recovers the classical Bayesian posterior. By training on synthetic data from a vast array of likelihood–prior pairs, TabPFN is encouraged to approximate this ideal behaviour across a highly diverse range of tasks.
@@
Why not other transformers? Large language models and other pretrained transformers often exhibit in-context learning: given a context $z_{1:i}$ and query $x_{i+1}$, they can produce a prediction for $y_{i+1}$. In this sense, they share the same basic mechanism as TabPFN. However, they lack the two further properties that make TabPFN especially well-suited as a predictive rule for martingale posteriors.
@@
First, they are not row-permutation invariant. Since such models rely on positional embeddings, their predictions change with the ordering of training examples, unless explicit averaging is performed across permutations. Second, they are not trained to approximate Bayesian PPD: their training objectives are tailored to next-token prediction in text rather than calibrated probabilistic prediction for supervised tabular tasks. These differences mean that, although generic transformers can perform in-context learning, their predictive distributions are not naturally aligned with the martingale posterior framework. By contrast, TabPFN combines in-context learning with built-in permutation invariance and Bayesian alignment, making it a more principled choice.
@@
Require : Observed data $z_{1:n}$, loss function $\ell(z,\theta)$, forward sampling depth NN, predictive rule $\mathrm{TabPFN}(\cdot\mid x_{i+1},z_{1:n})$, number of posterior samples $L$
@@
      Sample $y_{i+1}^{(l)}\leftarrow\mathrm{TabPFN}(\cdot\mid x_{i+1},z_{1:n})$;
@@
3.2 Validity of TabMGP.
A central question for any MGP construction is whether the chosen predictive rule leads to a well-defined limiting law $F_{\infty}$. While the framework traditionally points to conditions like the martingale property (Fong et al., 2023) or the almost conditionally identically distributed (a.c.i.d.) condition (Battiston and Cappello, 2025) as sufficient guarantees, verifying these for high-capacity models like TabPFN is currently impossible with existing tools in deep learning theory. Preliminary visual inspections of the predictive rule associated to TabPFN suggest there may be subtle deviations from these sufficient conditions (see Appendix C). However, we argue that such properties are sufficient rather than necessary for delivering useful epistemic uncertainty.
In the absence of formal theoretical guarantees, we rely on empirical diagnostics to verify the behaviour of TabMGP:
  • Path Stability (Scaled $L_{1}$ Convergence): The stability of the functional $\theta(F_{N})$ as the forward sampling depth $N$ increases is the primary diagnostic for the existence of a stable limit. We monitor this convergence by tracking the expected $L_{1}$-norm between the initial empirical risk minimiser $\theta(F_{n})$ and the terminal estimate $\theta(F_{N})$, scaled by the parameter dimension $p$: $\mathbb{E}_{F_{N}}[\frac{1}{p}\lVert\theta(F_{n})-\theta(F_{N})\rVert_{1}]$. If the transformer is internalising a consistent law, these trajectories should stabilise at a non-zero constant as $N$ grows. Systematic drift or divergent $L_{1}$ paths would indicate that there is no limiting measure.
  • Frequentist Coverage Assessment: We use frequentist coverage as an empirical check on the relationship between the transformer’s predictive rule and the population $F^{\star}$. A $(1-\alpha)$ credible set should contain the population risk minimiser $\theta(F^{\star})$ at a rate of at least $1-\alpha$.
  • Posterior Contraction: A valid epistemic uncertainty measure must represent uncertainty that vanishes with increasing data. One can verify this behaviour by observing the width of the posterior distribution of $\theta(F_{N})$ as the observed sample size $n$ increases. A sensible predictive rule should result in a distribution that narrows toward the population risk minimiser $\theta(F^{\star})$ as evidence accumulates, reflecting a reduction in epistemic uncertainty. See Figure 2 for an example.
Figure 2: Concentration of TabMGP. The black vertical line indicates the population risk minimiser $\theta(F^{\star})=(0,{\beta^{\star}}^{\top})$. See Appendix C for experimental details.
By focusing on these observed behaviours, practitioners can validate the outputs of TabMGP based on its empirical consistency, bypassing the currently insurmountable challenge of proving the transformer’s limiting properties.